OpenStudy (dreamie98):

A pyramid has a square base. If each edge of the pyramid is 4, what is the total surface area of the entire pyramid?

OpenStudy (varaug):

first you need to find the area of the square base. then use it to fill in the formula, [Base Area]+1/2 * perimeter * [slant length]

OpenStudy (dreamie98):

SA = 1/2ps + B SA = 1/2(4in x 4) (what is the height D:) + (4 x 4) SA = 1/2(16in)(WOT) + (16) I dont know what the slant height is :C

OpenStudy (dreamie98):

Is it also 4? @Varaug